FDC has decided to offer Unicorn Cookies.  We paid, a non-refundable, $120,000 for a marketing survey to help us understand food trends prior to settling in on Unicorn as the next new cookie option.  FDC thinks that the new cookie will generate $300,000 in incremental sales per year. Fixed costs will be $125,000 per year, and variable costs will be approximately 30% of sales (lots of food coloring).  The capital investment in the equipment needed to produce the new cookies will cost $200,000 and will be depreciated in a straight-line manner for the 4 years of the cookie’s life (if you think unicorn will really last that long, I seriously hope it is already over). Assume no salvage value Net working capital will not be affected by this project (you’re welcome). The firm has an average tax rate of 15% and a marginal tax rate of 21%.  The required rate of return on projects with similar risk is 9%.   

Lay it out before you start- you get credit for this…

                   

Time ZERO

Year 1-4

Cash Flow from Capital Investment

Cash Flow from Changes in Working Capital

Cash Flow from Ongoing Operations Annual Components:

         Revenue

                   Fixed

                   Variable

                Total Expenses

Depreciation

pretax

Taxes

After Tax Profit

Operating Cash flow

           

Please evaluate the cash flows for the project and calculate the NPV

Discount rate

         

         

Total NPV Calculation___________
Answer #1

Cash Flow from Capital Investment = 300000-125000 = 95000

Cash Flow from Changes in Working Capital = 120000+200000 =420000

Cash Flow from Ongoing Operations Annual Components:

Revenue = 95000

Fixed = 125000

Variable = 30% of Sales = 28500

Total Expenses = 153500

Depreciation = 50000

pretax = 311500

Taxes = 65415

After Tax Profit = 246085

Operating Cash flow = 268232

NPV = 88416.78